For 20 families we have the following frequency distribution of the number of different jobs the primary wage earner has held during the past five years. number of jobs (0X0 frequency 0 2 8 2 3 4 Find the relative frequency for X-1 jobs.
Answer #1

Answer

we know that relative frequency is defined as the individual frequency of the selected class or value divided by the total frequency

So, total frequency = 2+8+6+2+1+1 = 20

Corresponding frequency for X = 1 jobs is 8

Sp, relative frequency = (frequency of X = 1 jobs)/(total frequency) = (8/20) =0.40

So, required relative frequency is 0.40 (decimal form) or 0.40*100 = 40%